OpenStudy (jojo12g):

What value of k makes the statement true? –(k + 12) = 6 + (–12) A. –18 B. –6 C. 6 D. 18

OpenStudy (jojo12g):

b

TheSmartOne (thesmartone):

-(k+12) = -k - 12 -k + (-12) = 6 + (-12) can you tell what k is?

TheSmartOne (thesmartone):

correct :)

OpenStudy (jojo12g):

thanks

TheSmartOne (thesmartone):

anytime! :D